OpenStudy (anonymous):

what is the polar equation of a line passing through the pole with slope 1?

OpenStudy (anonymous):

HELP

zepdrix (zepdrix):

Hmm so we need to figure out this angle \(\Large\theta\), yes?

OpenStudy (anonymous):

we need to find the polar coordintes

OpenStudy (anonymous):

the polar equation

zepdrix (zepdrix):

Well since our line goes in both directions, it will include all values of r for the given angle. This polar equation will be of the form \(\Large\theta=b\). We need to figure out the angle.

zepdrix (zepdrix):

Since the slope of our line is 1, we can draw that relationship like this.\[\Large m\quad=\quad \frac{\Delta y}{\Delta x}\quad=\quad \frac{1}{1}\]